Ayasamrr:

The Sports Shop sells Nike running shoes for $82 a pair and Adidas running shoes for $95 a pair. One day, the Sports Shop sells 75 pairs of Nike and Adidas shoes totaling $6241 in sales. How many pairs of each shoe were sold?

mhanifa:

You need a system of equations. Let the number of Nike is x and number of Adidas is y. Can you build the two equations?

Ayasamrr:

i tried but im not sure

mhanifa:

Try the number of shoes, what is the sum?

Ayasamrr:

it is different then the total, can you solve it for me

mhanifa:

x + y = 75

mhanifa:

How much would be the cost of x Nike shoes and y Adidas shoes together?

mhanifa:

This is the second equation

Hajer:

1 attachment
Hajer:

this is the question can you solve it, because I'm confused *-*.

mhanifa:

The second equation is to determine the total cost. x Nike shoes cost 82x and y Adidas shoes cost 95y and their sum is 6241. How to show this as equation?

Hajer:

82x + 95y = 6241 , is it?

mhanifa:

@hajer wrote:
82x + 95y = 6241 , is it?
Yes Now you have both equations. Which method would you choose to solve the system: substitution or elimination?

Hajer:

substitution

Ayasamrr:

substitution if thats ok with u

mhanifa:

Ok, use the first equation. x + y = 75 We can put it as: x = 75 - y If you substitute x with 75 - y into the second equation, you will get an equation with one variable, y. Can you do that?

Hajer:

I will do it and send it to you to check if it is ok with you?

mhanifa:

@hajer wrote:
I will do it and send it to you to check if it is ok with you?
Its fine I'll wait

Hajer:

82(75 - y) + 95y = 6241 6150 - 82y + 95y = 6241 6150 + 13y = 6241 13y=91 y=7

mhanifa:

@hajer wrote:
82(75 - y) + 95y = 6241 6150 - 82y + 95y = 6241 6150 + 13y = 6241 13y=91 y=7
Excellent! Now you find the value of x easily

Ayasamrr:

thank you!!

Hajer:

Many thanks! ^-^

mhanifa:

You are welcome! And good luck with your assignment.